Solve by factoring and using the principle of zero products. 2 X - X = 20 e 1, 20 O 4,5 -4, -5 -4,5
Use the product rule to simplify the expression. 275 O 25 VTT 0 5VĪT 55
Answer #1

1. x^2 - x = 20

   => x^2 - x - 20 = 0

   => x^2 - 5x + 4x - 20 = 0

   => x(x - 5) + 4(x - 5) = 0

   => (x + 4)(x - 5) = 0

   => (x + 4) = 0, (x - 5) = 0

   => x = -4, x = 5

   Therefore,

   Answer: - 4, 5

2. \sqrt{275}

   = \sqrt{5 * 5 * 11}

   = \sqrt{5^2* 11}

   = 5\sqrt{11}